Finding analyatically limit of the function f(x)at x =a and hence verifying the continuinity function of that point
1
Expert's answer
2021-10-18T16:08:33-0400

for example: "f(x)=x^2"


"\displaystyle{\lim_{x\to a}}f(x)=\displaystyle{\lim_{x\to a}}x^2=a^2"


Since "\displaystyle{\lim_{x\to a}}f(x)=f(a)" , the function f(x) is continuous at x=a